OpenStudy (anonymous):

Show by hand how to find an anti-derivative to -t E^t^2

OpenStudy (solomonzelman):

u sub

OpenStudy (solomonzelman):

u=t²

jimthompson5910 (jim_thompson5910):

hint: u-subsitution u = t^2, so du = 2t*dt

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}f'(x)\times e^{f(x)}~dx}\) SUBSTITUTION: `u=f(x) du=f'(x) dx` \(\large\color{black}{\displaystyle\int\limits_{~}^{~}e^{u}~du=e^u+C=e^{f(x)}+C}\) (this is an abstract case for some differentiable f(x) )

OpenStudy (anonymous):

u =t^2 du = 2t dt -E^u du = ...

OpenStudy (solomonzelman):

well, not exactly

OpenStudy (anonymous):

I suppose (-1) can be pulled out

OpenStudy (nincompoop):

show by hand...

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}-t\times e^{t^2}~dt}\) \(\large\color{black}{\displaystyle-\int\limits_{~}^{~}t\times e^{t^2}~dt}\) \(\large\color{black}{\displaystyle u=t^2}\) \(\large\color{black}{\displaystyle du=2t~du~~~~\rightarrow~~~~\frac{1}{2}du=t~dt}\) \(\large\color{black}{\displaystyle-\frac{1}{2}\int\limits_{~}^{~} e^{u}~du}\)

OpenStudy (solomonzelman):

like that - same as you said, BUT you got the 1/2 there (why? I have wrote why in my post)

OpenStudy (anonymous):

and that 2 has to be removed I think.. du = 2t dt du/2 = 2t dt/2 du/2 = t dt ???

OpenStudy (anonymous):

thanks solomon...

OpenStudy (solomonzelman):

okay, can you now tell me what your antiderivative would be?

OpenStudy (solomonzelman):

(antiderivative of e^x is just e^x ... (well, +C) )

OpenStudy (anonymous):

-1/2 E^t^2

OpenStudy (solomonzelman):

yes, with +C

OpenStudy (anonymous):

oh yes.. +C

OpenStudy (solomonzelman):

One of my past professors took off most of the partial credit for +C. He explained that by saying that you gave only 1/∞ of all possible answers.

OpenStudy (anonymous):

thanks, nicely explained.

OpenStudy (anonymous):

my professors have been really good with that.

OpenStudy (solomonzelman):

Enjoy it any time I'm on. Although sometimes I will be gone. I wish good luck to you in all, Yes, that's the way! lets make it roll 0~0

OpenStudy (anonymous):

they had a grumble about how that happened to them.

OpenStudy (solomonzelman):

If you have further questions, then u know - whenever I'm.... (or, there are many other math peeps here) Just always watch out how you solve for du in your substitution ....

OpenStudy (anonymous):

actually if I had to use FTC with something like this equation.. for limits of [a,b] would they expect me to put +C on that too? say... -1/2 E^b^2 - (-1/2 E^a^2)

OpenStudy (solomonzelman):

wait, what does FTC stand for?

OpenStudy (anonymous):

fundamental theorem calculus

OpenStudy (solomonzelman):

oh.

OpenStudy (solomonzelman):

With limits of integration, +C ?

OpenStudy (anonymous):

I suppose it cancels out ...

OpenStudy (anonymous):

unless there's some kind of initial condition

OpenStudy (solomonzelman):

If you are integrating a definite integral, you don't have the +C, because when you are integrating a definite integral you are finding the (numerical) area under the curve over a specific interval. And you are not finding the family of functions. So they say I guess: \(\large\color{black}{\displaystyle\int\limits_{a}^{b}f(x)~dx=\left( f(x)+C\right){\LARGE |}^{x=b}_{x=a}}\) (where F'=f)

OpenStudy (solomonzelman):

and yes C cancel's out, but it shouldn't be there in a first place.

OpenStudy (solomonzelman):

+C is a family of antiderivative functions F(x) (and this family of F(x)+C is a set of functions that has a derivative of f) when finding area under the corve over some interval I , that is not a family of functions you are finding - that is an area calculation.

OpenStudy (solomonzelman):

Not that the result would differ if you have +C there, since it will cancel, but as far as the concept goes +C, again, should not be there)

OpenStudy (anonymous):

gotcha.. makes sense

OpenStudy (solomonzelman):

Alrighty:)

OpenStudy (anonymous):

laterzs

OpenStudy (anonymous):

thanks again

OpenStudy (solomonzelman):

Anytime